OpenStudy (anonymous):

Find the surface area of a square pyramid, where the side of the square is 16 in and the height of the pyramid h=6 in

OpenStudy (anonymous):

Surface square pyramid = 4* surface triangle + surface square. Surface triangle = 0.5*b*h b=16 in h=sqrt((0.5*16)²+6²)=sqrt(8²+6²)=sqrt(100)=10 in Surface triangle = 0.5*16*10=80 in² Surface square = side * side Surface square = 16*16=256 in Surface square pyramid = 4 * 80 + 256 = 320 + 256 = 576 in²